An investor invested a total of $2,500.00 in two mutual funds. one fund earned 9% of the profit while the other earned a 6% profit. If the investors total profit was $165.00 how much was invested in each mutual fund?
Found 2 solutions by mananth, amalm06:
Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
2500
Interest----- 165.00

Part I 9.00% per annum ---x *
Part II 6.00% per annum ---y
Total investment
x + 1 y= 2500 -------------1
Interest on both investments
9.00% x + 6.00% y= 165
Multiply by 100
9 x + 6 y= 16500.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -22500.00
Add this to (2)
0 x -3 y= -6000
divide by -3
y = 2000 - x)
Part I 9.00% $ 500
Part II 6.00% $ 2000

CHECK
500 --------- 9.00% ------- 45.00
2000 ------------- 6.00% ------- 120.00
Total -------------------- 165.00
11.6
m.ananth@hotmail.ca 1520

Answer by amalm06(224)

Solve with one equation:
0.09x+(2500-x)(0.06)=165
0.09x+150-0.06x=165
0.03x=15
x=$500
$500 at 9% (Answer), $2000 at 6% (Answer)